irm L had net assets at the end of the year of $390,000 .The only transactions affecting stockholders’ equity during the year were net income of $60,000 and dividends of $30,000. Required: Calculate Firm L’s average stockholders’ equity and return on equity (ROE).
Answer #1

Solution:

Beginning SE = Ending SE - Net income + Dividends = 390,000 - 60,000 + 30,000 = 360,000

Average stockholders’ equity = (Beginning SE + Ending SE) / 2 = (390,000 + 360,000) / 2 = 375,000

ROE = Net income / Average stockholders’ equity = 60,000 / 375,000 = 0.16 or 16%